# MIST: Mutual Information Estimation via Supervised Training

### v1

@@ 1 Introduction
Mutual information (MI) is a measure of nonlinear statistical dependence between two random variables, quantifying how much knowing one variable reduces the uncertainty about the other (shannon1948mathematical; mackay2003information). For two continuous random variables $X$ and $Y$, the MI is written in terms of their joint distribution $P_{(X,Y)}$ and marginal distributions $P_{X}$ and $P_{Y}$:
@@
Mutual information is one of the most widely applied information measures in data science, finding applications in feature selection (peng2005feature_selection; kwak2002_feature_selection), causality (butte1999_causality_genomics), representation learning (chen2016infogan; higgins2018towards_distentagled_representations; tishby2015_info_bottleneck; zhao2018information_autoencoders), reinforcement learning (oord2018_cpc; pathak2017curiosity), and generative modeling (alemi2016deep_variation_info_bottleneck; alemi2018gilbo; huang2020evaluating_lossy_compression).
Calculating MI requires access to the joint and marginal distributions, which are rarely known in practice. Consequently, most approaches infer MI from finite samples drawn from these distributions. Existing estimators fall into two main families. (1) Density estimators approximate the underlying densities, for example, using kernel density estimation, $k$-nearest neighbors, or variational autoencoders. MI can then be computed by substituting these estimates into one of MI’s definitions (such as Equation 1). Prominent density estimators include KSG (ksg) and MINDE (minde). (2) Density relationship estimators instead estimate the relationship between the joint and marginal distributions, often using variational lower bounds derived from the Donsker–Varadhan representation of the Kullback–Leibler divergence (see Appendix A.1 for these definitions). Notable examples include MINE (mine), NWJ (nwj), and SMILE (MINE_exp_variance_consistency_tests).
@@
In both families of approaches, neural methods have been developed to improve the estimator’s capacity to capture complex density functions or density relationships. Broadly, these neural methods have achieved better performance on moderately complex distributions and high-MI settings. However, both families of approaches struggle in challenging settings that are common in application, including high-dimension, limited samples, complex distributions, and high MI (bmi_benchmark; mcallester2020formal_limitations). Prominent examples of these settings include applications in genomics (klein2015droplet_genomics), neuroscience (ganguli2012compressed_neuroscience), astronomy (pandey2017much_astronomy), and machine learning (goldfeld2018estimating_info_flow). While previous benchmarking efforts have expanded the set of distribution families considered, MI estimators have rarely been empirically evaluated under these challenging conditions individually, and even more rarely in combination. Most empirical evaluations have studied estimation for MI below 1, for fewer than 5 distributions, with a minimum sample size above 10,000, and for higher dimensions limited to normal distributions. See Appendix A.3 for an overview of the empirical evaluation scale for existing methods.
@@
In this work, we propose a fully data-driven, empirical approach to designing mutual information estimators. Instead of building estimators that compute MI by approximating density functions or their relation, we learn to predict MI directly from data. We parameterize a mutual information estimator as a neural network and train such a network end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. Our proposed model, MIST (Mutual Information estimation via Supervised Training), handles arbitrary sample sizes and dimensions, and features built-in uncertainty estimates. This meta-learning approach takes inspiration from amortized inference procedures such as simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a), and meta-statistical learning (peyrard2025meta).
@@
Our work focuses on the setting for MI estimators that was underexplored in previous work. Empirically, we conduct a large-scale evaluation that systematically explores the performance of our estimator in the most difficult regimes: with fewer samples (10-500), in higher dimensions (2-32), and across a wider range of MI values (0-40). We find that our learned estimator outperforms existing methods across these conditions, especially in low-sample regimes. We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that our models require roughly half as many samples on average to provide reliable MI estimates as the best baseline. The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives, while inference is orders of magnitude faster than existing neural methods. Finally, the learned estimators generalize to unseen distributions and sample sizes.
@@
Beyond empirical performance, this meta-statistical framework establishes a new paradigm for constructing information-theoretic estimators. Because the estimator is fully differentiable and trainable, it can be embedded in larger learning pipelines, fine-tuned to specific data modalities, and systematically improved through curated meta-datasets. Leveraging MI’s invariance under invertible transformations, synthetic training distributions can, in the future, be adapted to arbitrary domains and modalities via normalizing flows (papamakarios2021normalizing_flows), paving the way for general-purpose, learned estimators of mutual information.
We provide an open-source library [^1] https://github.com/grgera/mist for training and evaluating meta-learned MI estimators, and include a detailed discussion of these opportunities and future work in Section 5.
@@ 2 Related Work
Mutual information estimation is a foundational challenge across data science tasks, for which a wide variety of estimation methods have been developed. Although these approaches vary in their mathematical foundations and computational strategies, they can be grouped into two main families: density estimators and density ratio estimators.
@@
Density estimators compute mutual information estimates by performing intermediate estimates of density functions. The most straightforward approach is to use MI’s analytical definition, which requires estimates of the joint and marginal density functions. Histogram MI estimators approximate these density functions by partitioning the support of each variable into a set of bins and counting the frequencies in each (histogram_binning). Histogram estimators suffer from very poor scaling and are highly sensitive to the choice of bins, leading to research on optimizing the bin selection (adaptive_binning). Kernel density estimators (KDE; kde; textbook_kde) were later developed to help avoid this sensitivity by replacing fixed bins with kernel functions centered at each sample point.
original_kl_density_estimator designed a method for entropy estimation using neighborhood distances, an approach which was later refined with weighted variants (weighted_kl_density_estimator). The Kraskov-Stögbauer-Grassberger estimator (KSG; ksg) extended this k-NN framework to mutual information estimation using MI’s expression in terms of conditional entropy (see Appendix A.1). KSG is one of the most broadly used MI estimators, especially for low-sample settings, and serves as a primary benchmark comparison in this work.
@@
Several approaches have explored the use of deep generative models for density estimation. One line of work uses normalizing flows to reconstruct density functions directly. However, these methods suffer from poor estimates even for data with simple structure (MINE_exp_variance_consistency_tests; dine_gaussian). The Barber-Agakov bound (BA; barber2004algorithm_BA_Bound) provides a variational lower bound on MI by introducing a variational approximation $q(x|y)$ to the true conditional $p(x|y)$. Several works have applied this bound using conditional generative models with different choices for the marginal distribution (alemi2016deep_variation_info_bottleneck; chalk2016relevant_BA_bound_application; kolchinsky2019nonlinear). Later work used VAEs to learn the joint and marginal density functions directly (MINE_exp_variance_consistency_tests). Most recently, MINDE (minde) trains conditional or joint diffusion models to learn score functions, then applies the Girsanov theorem to compute KL-divergence as the expected difference between joint and marginal scores. VCE (chen2025neural_vector_copulas) is a concurrent work to ours, and writes MI as the negative differential entropy of the vector copula density. Their approach uses flow models to estimate the marginal densities, then estimates the vector copula density using maximum likelihood estimation.
@@
Another line of density estimators has focused on building MI estimators for complex settings by learning methods to simplify the distributions before applying classical estimators or computing MI directly. MIENF (mi_through_normalizing_flows) learns a pair of normalizing flows that apply MI-preserving transformations to the joint distribution, enabling closed-form MI calculation. Latent-MI (LMI; lmi_estimator) trains a network to compress data with minimal mutual information loss, then applies KSG to the low-dimensional embeddings. These methods have improved abilities to handle high-dimensional data, but have large sample and computational requirements, and rely on the existence of simple or low-dimensional structure in high-dimensional data.
@@
Density ratio estimators compute mutual information by estimating the relationship between the joint density $P(X,Y)$ and the product of marginal densities $P_{X}P_{Y}$. Since mutual information can be equivalently expressed as the Kullback-Leibler divergence $D_{KL}(P(X,Y)\|P_{X}P_{Y})$ (donsker_varadhan_kl_div_mi_definition), many density ratio methods leverage variational bounds on KL divergence. Several methods have been developed by defining discriminative tasks that implicitly predict this density ratio. MINE (mine) trains a neural network critic to differentiate joint and marginal samples, optimizing the critic to maximize the Donsker-Varadhan (DV) lower bound. The Nguyen-Wainwright-Jordan estimator (NWJ; nwj) trains a critic model similarly, but uses an alternative variational bound which offers lower variance at the cost of a looser bound. InfoNCE (oord2018_cpc) trains a classifier critic to identify true joint samples among negative samples drawn from the marginals. SMILE (MINE_exp_variance_consistency_tests) combines the InfoNCE and NWJ bounds through a clip operation, interpolating between their respective bias-variance profiles to achieve more flexible trade-offs. The limitations of variational estimators are well-studied in mcallester2020formal_limitations, MINE_exp_variance_consistency_tests, and poole_variation_bounds_survey, and include high variance estimates, high sample requirements, and poor estimates for high MI settings. We refer the reader to poole_variation_bounds_survey for a detailed review of MI variational bounds.
Evaluation of MI Estimators.
Evaluating MI estimators is itself challenging: it requires distributions with known ground-truth MI and large sample sizes (required by most methods). Existing work has therefore primarily focused on theoretical guarantees of low bias, consistency with increasing samples, and the tightness of variational bounds (mine; nwj; MINE_exp_variance_consistency_tests). Empirical evaluations have been limited in scope. Early methods, such as binning and kernel density estimators, were tested on simple time-series data like sine waves (kde). Later neural methods began evaluation on a smaller set of synthetic distributions: MINE and SMILE were evaluated on 3–4 distributions, including high-dimensional Gaussians up to 20D (mine; MINE_exp_variance_consistency_tests). The majority of methods have been evaluated exclusively on MI values below 1, often on Gaussian or near-Gaussian distributions (see Table 3). Methods that consider higher dimensionalities do so largely by scaling low-dimensional distributions (assuming low-dimensional latent structure), or test only on high-dimensional Gaussians (lmi_estimator; minde).
Recently, the BMI benchmark (bmi_benchmark) applied 20 invertible transformations to normal and Student’s t-distributions to create a broader set of distribution families with unique challenges, such as sparsity and long tails. This evaluation on more complex distributions highlighted the limitations of many existing estimators, leading authors to argue that multivariate normal distributions provided a overly optimistic view of estimator performance. BMI has been used by some of the most recent estimators, such as MINDE. We provide a description of our datasets in Appendix B, and an overview of existing evaluation scopes in Appendix A.3.
@@
Our work aligns with the broader research direction on learning inference procedures. A relevant related direction concerns neural processes (garnelo2018neural; garnelo2018conditional; kim2019attentive; Gordon:2020; markou2022practical; huang2023practical; BruinsmaMRFAVBH23), which can predict latent variables of interest from datasets (chang2025amortized) by leveraging transformers (pmlr-v162-nguyen22b) and Deep Sets (NIPS2017_f22e4747) to enforce permutation invariance (JMLR:v21:19-322). A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (muller2022transformers) and optimization tasks (pmlr-v202-muller23a). Additionally, there is growing interest in using trained models to assist in statistical inference tasks (angelopoulos2023predictionpoweredinference) and optimization (NIPS2017_addfa9b7; NEURIPS2020_f52db9f7; NEURIPS2021_56c3b2c6; amos2023tutorialamortizedoptimization) like simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a) and the learning of frequentist statistical estimators (peyrard2025meta).
@@ 3 Learning Framework
The methods available today decompose MI estimation into intermediate tasks of density function or density ratio estimation. To estimate a density function (or ratio) beyond very simple settings, these methods require large amounts of samples as part of inference. Even with access to a high number of samples, density estimation itself is a notoriously difficult and unsolved problem, particularly for high-dimensional or complex distributions (density_estimation). As a result, existing MI estimators still fail in many realistic data settings. They scale poorly to high-dimensional data and complex distributions, are unstable for low-sample settings, and require expensive inference procedures. An additional consequence is that studying and improving MI estimators is challenging, with new methods often resulting from theoretical discoveries.
Our work offers an alternative philosophy for designing mutual information estimators. Instead of building estimators that compute MI with a pre-specified algorithm (e.g., approximating density functions or their ratios), we propose to predict MI directly from data. We take a fully empirical perspective, asking: Can a model learn a generalizable algorithm for predicting mutual information directly from samples?
To learn this algorithm, we parameterize a mutual information estimator as a neural network that we call MIST (Mutual Information estimation from Supervised Training). MIST takes as input a set of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $P_{XY}$, and outputs an estimate of $I(X;Y)$. The model is trained end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. We have reflected our framework in Figure 1 and described the meta-dataset design, architecture, and training choices below.
@@
Training our meta-learned estimators in a supervised manner requires distributions with known MI. We generate these distributions using BMI, a library designed to create complex benchmarks for MI estimation. BMI exploits the invariance of mutual information to homeomorphisms, applying invertible transformations to simple base distributions to produce more complex and challenging distributions. We partition the base distribution families into separate train and test sets, that is, the meta-distribution used for training has a different support than the one used for testing. For each base distribution, we apply BMI’s suite of invertible transformations and generate variants across dimensions ranging from $2$ to $32$, yielding two distribution sets $\Gamma_{\text{train}}$ and $\Gamma_{\text{test}}$. From these distributions, we construct a training meta-dataset $\mathcal{M}_{\text{train}}$ and two test meta-datasets differing only in scale: $\mathcal{M}_{\text{test}}$, and $\mathcal{M}_{\text{test-extended}}$. These test meta-datasets contain a mix of distribution families found in $\Gamma_{\text{train}}$ (subsequently referred to as in-meta-distribution or IMD) or unseen at training time (out-of-meta-distribution or OoMD). Since MI is a function of two random variables, each meta-datapoint in our meta-datasets consists of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $\gamma\in\Gamma$, along with the known mutual information $I_{\gamma}(X;Y)$ as the label. The value of $n$ varies between $10$ and $500$ across each of our train and test meta-datasets, thus focusing on low-sample regimes. The resulting training meta-dataset $\mathcal{M}_{\text{train}}$ contains 16 base distribution families and 625,000 meta-datapoints. $\mathcal{M}_{\text{test-extended}}$ contains 806,000 meta-datapoints from 13 distribution families, of which 6 are shared with $\mathcal{M}_{\text{train}}$. Meanwhile, $\mathcal{M}_{\text{test}}$ has a reduced size of 2,340 meta-datapoints, created to allow the testing of slower baselines that would take prohibitively long to run on the extended test set. The test sets are sampled to uniformly cover all sample sizes, dimensions, and base distribution families. The contents and generation of the meta-datasets are described in more detail in Appendix B.
@@
Building an estimator using our framework presents three main architectural challenges. First, the model must process datasets of varying size while remaining invariant to sample ordering. The Transformer architecture is naturally permutation-invariant and has been a common choice to ingest datasets as inputs (transformer_invariance). The SetTransformer (set_transformer) architecture introduced the ISAB block to reduce the quadratic cost of attention by performing attention on a fixed number of learned inducing points. Finally, the SetTransformer++ architecture (set_transformer_plus_plus) further improves upon the original SetTransformer by introducing set normalization and additional residual connections that enable more stable training with deeper models. We therefore use SetTransformer++ as the basis of our architecture. Second, our estimator must process samples of varying dimensions. To address this, we introduce an additional attention block that operates over the dimension axis; this time, permutation invariance is not required. This row-wide attention also operates with a learned fixed inducing point, which then serves as a pooling mechanism into a fixed row dimension. Third, unlike many supervised learning tasks, mutual information is an unbounded measure. This is a long-standing challenge for MI estimators, which typically underestimate MI and fail in high-MI settings. We explored several normalization and multi-task prediction strategies (e.g., predicting normalized conditional entropies) to address this problem when designing our model, and provide our results in Appendix C.1. We found that the simplest solution of directly predicting MI produced the best estimators, provided the meta-dataset includes a sufficient range of MI labels.

### v3

@@ 1 Introduction
Mutual information (MI) is a measure of nonlinear statistical dependence between two random variables, quantifying how much knowing one variable reduces the uncertainty about the other (Shannon, 1948; MacKay, 2003). For two continuous random variables $X$ and $Y$, the MI is written in terms of their joint distribution $P_{(X,Y)}$ and marginal distributions $P_{X}$ and $P_{Y}$:
@@
Mutual information is one of the most widely applied information measures in data science, finding applications in feature selection (Peng et al., 2005; Kwak and Choi, 2002), causality (Butte and Kohane, 1999), representation learning (Chen et al., 2016; Higgins et al., 2018; Tishby and Zaslavsky, 2015; Zhao et al., 2018), reinforcement learning (Oord et al., 2018; Pathak et al., 2017), and generative modeling (Alemi et al., 2016; Alemi and Fischer, 2018; Huang et al., 2020).
@@
Calculating MI requires access to the joint and marginal distributions, which are rarely known in practice. Consequently, most approaches infer MI from finite samples drawn from these distributions. Existing estimators fall into two main families: density estimators and density ratio estimators. Density estimators approximate the underlying densities, for example, using kernel density estimation, $k$-nearest neighbors, or variational autoencoders. MI can then be computed by substituting these estimates into one of MI’s definitions (such as Equation 1). Prominent density estimators include KSG (Kraskov et al., 2004) and MINDE (Franzese et al., 2023). Density ratio estimators instead estimate the relationship between the joint and marginal distributions, often using variational lower bounds. Notable examples include MINE (Belghazi et al., 2018), NWJ (Nguyen et al., 2010), and SMILE (Song and Ermon, 2019). In both families of approaches, neural methods have been developed to improve the estimator’s capacity to capture complex density functions or density relationships. Broadly, these neural methods have achieved better performance on moderately complex distributions and high-MI settings. However, both families of approaches struggle in challenging settings that are common in practice: high dimensionality, limited samples, complex distributions, and high mutual information (Czyż et al., 2023; McAllester and Stratos, 2020). Examples of these settings include applications in genomics (Klein et al., 2015), neuroscience (Ganguli and Sompolinsky, 2012), astronomy (Pandey and Sarkar, 2017), and machine learning (Goldfeld et al., 2018). While previous benchmarking efforts have expanded the set of distribution families considered, MI estimators have rarely been empirically evaluated under these challenging conditions individually, and even more rarely in combination. Most empirical evaluations have studied estimation for fewer than 5 distributions, with a minimum sample size of 10,000 or more, and in higher dimensions primarily for normal distributions (Appendix A.3). Existing work has instead primarily focused on theoretical guarantees of low bias, consistency with increasing samples, and the tightness of variational bounds (Belghazi et al., 2018; Nguyen et al., 2010; Song and Ermon, 2019).
@@
In this work, we propose a fully data-driven, empirical approach to designing mutual information estimators. Instead of building estimators that compute MI by approximating density functions or their relation, we learn to predict MI directly from data. We parameterize a mutual information estimator as a neural network and train such a network end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. Our proposed model, MIST (Mutual Information estimation via Supervised Training), handles arbitrary sample sizes and dimensions, and features built-in uncertainty estimates. This meta-learning approach takes inspiration from amortized inference procedures such as simulation-based inference (SBI; Papamakarios et al., 2019; Cranmer et al., 2020), amortized Bayesian inference (ABI; Gonçalves et al., 2020; Elsemüller et al., 2024; Radev et al., 2023; Avecilla et al., 2022; Gloeckler et al., 2024), and meta-statistical learning (Peyrard and Cho, 2025).
@@
We focus on empirical settings underexplored in previous work but common in practice: fewer samples (10-500), higher dimensions (2-32), and a wider range of MI values (0-40) across 23 distribution types. We find that MIST substantially outperforms existing methods across these conditions, especially in low-sample regimes. We provide a detailed study of sample requirements under increasing dimensionality, finding that MIST requires approximately half as many samples on average to provide reliable MI estimates compared to the next best estimator. The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives, while inference is orders of magnitude faster than existing neural methods. Finally, MIST generalizes effectively to unseen distributions and sample sizes.
@@
Beyond empirical performance, this meta-statistical framework establishes a new paradigm for constructing information-theoretic estimators. Because the estimator is fully differentiable and trainable, it can be embedded in larger learning pipelines, fine-tuned to specific data modalities, and systematically improved through curated meta-datasets. Leveraging MI’s invariance under invertible transformations, synthetic training distributions can, in the future, be adapted to arbitrary domains and modalities via normalizing flows (Papamakarios et al., 2021), paving the way for general-purpose, learned estimators of mutual information. We provide an open-source library [^1] https://github.com/grgera/mist for training and evaluating meta-learned MI estimators, and include a detailed discussion of these opportunities and future work in Section 5.
@@ 2 Related Work
Mutual information estimation is a foundational challenge across data science tasks, for which a wide variety of estimation methods have been developed. Although these approaches vary in their mathematical foundations and computational strategies, they can be grouped into two main families: density estimators and density ratio estimators. See Appendix A.2 for a detailed discussion of previous estimators and our choices in categorizing them.
@@
Density estimators compute mutual information estimates by performing intermediate approximations of density functions. Histogram MI estimators approximate the joint and marginal density functions in 1 by partitioning the support of each variable into a set of bins and counting the frequencies in each (Fraser and Swinney, 1986). Histogram estimators suffer from very poor scaling and are highly sensitive to the choice of bins, leading to research on optimizing the bin selection (Darbellay and Vajda, 1999). Kernel density estimators (KDE; Moon et al., 1995; Silverman, 2018) were later developed to help avoid this sensitivity by replacing fixed bins with kernel functions centered at each sample point. Kozachenko (1987) designed a method for entropy estimation using neighborhood distances, an approach which was later refined with weighted variants (Berrett et al., 2019). The Kraskov-Stögbauer-Grassberger estimator (KSG; Kraskov et al., 2004) extended this k-nearest neighbor framework to mutual information estimation using MI’s expression in terms of conditional entropy (see Appendix A.1). KSG is one of the most broadly used MI estimators, especially for low-sample settings, and serves as a primary benchmark comparison in this work. Several approaches have explored the use of deep generative models for density estimation.
@@
One line of work uses VAEs or normalizing flows to reconstruct the joint and marginal density functions directly, an approach which suffers from high sample requirements and a limited ability to capture complex structure (Song and Ermon, 2019; Duong and Nguyen, 2023). The Barber-Agakov bound (BA; Barber and Agakov, 2004) provides a variational lower bound on MI by introducing a variational approximation $q(x|y)$ to the true conditional $p(x|y)$. Several works have applied this bound using conditional generative models with different choices for the marginal distribution (Alemi et al., 2016; Chalk et al., 2016; Kolchinsky et al., 2019). More recently, MINDE (Franzese et al., 2023) trains conditional or joint diffusion models to learn score functions, then applies the Girsanov theorem to compute KL-divergence as the expected difference between joint and marginal scores. VCE (Chen et al., 2025) is a concurrent work that expresses MI as the negative differential entropy of the vector copula density, which they estimate by first using normalizing flows to model the marginal densities, then applying maximum likelihood estimation to the copula.
@@
Another line of density estimators has focused on building MI estimators for complex settings by learning methods to simplify the distributions before applying classical estimators or computing MI directly. MIENF (Butakov et al., 2024) learns a pair of normalizing flows that apply MI-preserving transformations to the joint distribution, enabling closed-form MI calculation. Latent-MI (LMI; Gowri et al., 2024) trains a network to compress data with minimal mutual information loss, then applies KSG to the low-dimensional embeddings. These methods have improved abilities to handle high-dimensional data, but have large sample and computational requirements, and rely on the existence of simple or low-dimensional structure in high-dimensional data.
@@
Density ratio estimators compute mutual information by estimating the relationship between the joint density $P(X,Y)$ and the product of marginal densities $P_{X}P_{Y}$. Since mutual information can be equivalently expressed as the Kullback-Leibler divergence $D_{KL}(P(X,Y)\|P_{X}P_{Y})$ (Donsker and Varadhan, 1975), many density ratio methods leverage variational bounds on KL divergence. Several methods have been developed by defining discriminative tasks that implicitly predict this density ratio. MINE (Belghazi et al., 2018) trains a neural network critic to differentiate joint and marginal samples, optimizing the critic to maximize the Donsker-Varadhan (DV) lower bound. The Nguyen-Wainwright-Jordan estimator (NWJ; Nguyen et al., 2010) trains a critic model similarly, but uses an alternative variational bound which offers lower variance at the cost of a looser bound. InfoNCE (Oord et al., 2018) trains a critic to identify true joint samples among negative samples drawn from the marginals. SMILE (Song and Ermon, 2019) combines the InfoNCE and NWJ bounds through a clip operation, interpolating between their respective bias-variance profiles to achieve more flexible trade-offs. The limitations of variational estimators are well-studied in McAllester and Stratos (2020), Song and Ermon (2019), and Poole et al. (2019), and include high variance estimates, high sample requirements, and poor estimates for high MI settings. We refer the reader to Poole et al. (2019) for a detailed review of MI variational bounds.
FMMI (Butakov et al., 2025) learns a coupling transform to map between the joint distribution and marginal product, using the expected divergence of the velocity field as the MI estimate. InfoNet (Hu et al., 2024) trains a neural network to predict the output of a trained MINE critic, removing the need to train a critic on each new set of samples. Like our work, this approach achieves substantial test-time efficiency gains through large-scale pretraining. However, our works differ in the learning objective: InfoNet learns to emulate an algorithm (MINE), whereas MIST learns an algorithm that allows it to predict the MI directly from samples.
@@
Our work aligns with the broader research direction on learning inference procedures. A relevant related direction concerns neural processes (Garnelo et al., 2018b; a; Kim et al., 2019; Gordon et al., 2020; Markou et al., 2022; Huang et al., 2023; Bruinsma et al., 2023), which can predict latent variables of interest from datasets (Chang et al., 2025) by leveraging transformers (Nguyen and Grover, 2022) and Deep Sets (Zaheer et al., 2017) to enforce permutation invariance (Bloem-Reddy and Teh, 2020). A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (Müller et al., 2022) and optimization tasks (Müller et al., 2023). Additionally, there is growing interest in using trained models to assist in statistical inference tasks (Angelopoulos et al., 2023) and optimization (Lueckmann et al., 2017; Liu et al., 2020; Simpson et al., 2021; Amos, 2023) like simulation-based inference (SBI; Papamakarios et al., 2019; Cranmer et al., 2020), amortized Bayesian inference (ABI; Gonçalves et al., 2020; Elsemüller et al., 2024; Radev et al., 2023; Avecilla et al., 2022; Gloeckler et al., 2024) and the learning of frequentist statistical estimators (Peyrard and Cho, 2025).
@@ 3 Learning Framework
The methods available today decompose MI estimation into intermediate tasks of density function or density ratio estimation. To estimate a density function (or ratio) beyond simple settings, existing methods require large amounts of samples as part of inference. Even with access to a high number of samples, density estimation itself is a notoriously difficult and unsolved problem, particularly for high-dimensional or complex distributions (Mészáros et al., 2024). As a result, the MI estimators available today still frequently fail in many realistic data settings. They scale poorly to high-dimensional data and complex distributions, are unstable for low-sample settings, and require expensive inference procedures. An additional consequence is that studying and improving MI estimators is challenging, with new methods often resulting from theoretical discoveries.
Our work offers an alternative philosophy for designing mutual information estimators. Instead of building estimators that compute MI with a pre-specified algorithm, such as approximating density functions or their ratios, we propose to predict MI directly from data. We take a fully empirical perspective, asking: Can a model learn a generalizable algorithm for predicting mutual information directly from samples? To learn this algorithm, we parameterize a mutual information estimator as a neural network that we call MIST (Mutual Information estimation via Supervised Training). MIST takes as input a set of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $P_{XY}$, and outputs an estimate of $I(X;Y)$. The model is trained end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. We provide a visualization of our framework in Figure 1 and describe the meta-dataset design, architecture, and training choices below.
@@
Training our meta-learned estimators in a supervised manner requires distributions with known MI. We generate these distributions using BMI (Czyż et al., 2023), a library designed to create complex benchmarks for MI estimation. BMI exploits the invariance of mutual information to homeomorphisms, applying invertible transformations to simple base distributions to produce more complex and challenging distributions. We partition the base distribution families into separate train and test sets, that is, the meta-distribution used for training has a different support than the one used for testing. For each base distribution, we apply BMI’s suite of invertible transformations and generate variants across dimensions ranging from $2$ to $32$, yielding two distribution sets $\Gamma_{\text{train}}$ and $\Gamma_{\text{test}}$.
From these distributions, we construct a training meta-dataset $\mathcal{M}_{\text{train}}$ and two test meta-datasets differing only in scale: $\mathcal{M}_{\text{test}}$, and $\mathcal{M}_{\text{test-extended}}$. Both scales of test meta-datasets contain an in-meta-distribution (IMD) subset, in which the distribution family was seen in $\Gamma_{\text{train}}$, and an out-of-meta-distribution (OoMD) subset. Each meta-datapoint in our meta-datasets consists of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $\gamma\in\Gamma$, along with the known mutual information $I_{\gamma}(X;Y)$ as the label. The value of $n$ varies between $10$ and $500$ across each of our train and test meta-datasets. The resulting training meta-dataset $\mathcal{M}_{\text{train}}$ contains 16 base distribution families and 625,000 meta-datapoints. $\mathcal{M}_{\text{test-extended}}$ contains 806,000 meta-datapoints from 13 distribution families, of which 6 are shared with $\mathcal{M}_{\text{train}}$. Meanwhile, $\mathcal{M}_{\text{test}}$ has a reduced size of 2,340 meta-datapoints, created to allow the testing of slower baselines that would take prohibitively long to run on the extended test set. The test sets are sampled to uniformly cover all sample sizes, dimensions, and base distribution families. The contents and generation of the meta-datasets are described in more detail in Appendix B.
@@
Building an estimator using our framework presents three main architectural challenges. First, the model must process datasets of varying size while remaining invariant to sample ordering. The Transformer architecture is naturally permutation-invariant and has been a common choice to ingest datasets as inputs (Xu et al., 2024). The SetTransformer (Lee et al., 2019) architecture introduced the ISAB block to reduce the quadratic cost of attention by performing attention on a fixed number of learned inducing points. Finally, the SetTransformer++ architecture (Zhang et al., 2022) further improves upon the original SetTransformer by introducing set normalization and additional residual connections that enable more stable training with deeper models. We therefore use SetTransformer++ as the basis of our architecture.
Second, our estimator must process samples of varying dimensions. To address this, we introduce an additional attention block that operates over the dimension axis; this time, permutation invariance is not required. This row-wide attention also operates with a learned fixed inducing point, which then serves as a pooling mechanism into a fixed row dimension. Third, unlike many supervised learning tasks, mutual information is an unbounded measure. This is a long-standing challenge for MI estimators, which typically underestimate MI and fail in high-MI settings. We explored several normalization and multi-task prediction strategies to address this problem when designing our model, and provide our results in Appendix C.1. We found that the simplest solution of directly predicting MI produced the best estimators, provided the meta-dataset includes a sufficient range of MI labels.